After touchdown a fighter jet passes a marker on the runway at an instantaneous speed of 100 m/s and constant negative acceleration of -15 m/s2. After slowing down to 25 m/s, the pilot suddenly goes to full power to provide a new constant acceleration of 34 m/s2 for a "touch and go" takeoff. How far from the original runway marker does the plane achieve its takeoff speed of 147 m/s?
Answer #1

during deceleration


v1x^2 - vox^2 = 2*a1x*x1

25^2 - 100^2 = 2*(-15)*x1

x1 = 312.5 m


durin acceleration


v2x^2 - v1x^2 =2*a2x*x2


147^2 - 25^2 = 2*34*x2


x2 = 308.6 m

total distance from the original runway = x1 + x2 = 621.1 m